The first recorded observation of hydrogen gas was recorded by the famous alchemist paracelsus when he added iron to sulfuric acid. Calculate the mass of hydrogen gas produced when paracelsus dissolved 10.0g of iron in excess sulfuric acid

2Fe + 3H2SO4 --> Fe2SO4 + 3H2
1
Expert's answer
2020-06-08T15:04:49-0400

2Fe + 3H2SO4 --> Fe2(SO4)3 + 3H2

n(Fe) = 10.0g / 55.8g/mol = 0.18 mol

According to the equation, n(H2) = 3/2 x n(Fe) = 0.27 mol

m(H2) = 0.27mol x 2.0g/mol = 0.54 g

Answer: 0.54 g